OpenStudy (anonymous):

The function f is defined by f(x)=5-3sin2x for 0

OpenStudy (experimentx):

,min value is when 3sin(2x) = 3, 2x = pi/2, x=> pi/4, min value = 5 - 3 = 2 max value is when 3sin(2x) = -3, 2x=3pi/2, x = 3pi/4, max value = 5 - (-3) = 8 range = (2,8)

OpenStudy (anonymous):

So simply find the derivative and set it equal to zero. The values you find plug them back in to the original and that is your range?

OpenStudy (anonymous):